# Human-AI Collaboration in Cloud Security: Cognitive Hierarchy-Driven Deep Reinforcement Learning

### v1

@@ Abstract.
Given the complexity of multi-tenant cloud environments and the growing need for real-time threat mitigation, Security Operations Centers (SOCs) must integrate AI-driven adaptive defense mechanisms to counter Advanced Persistent Threats (APTs) effectively. However, SOC analysts face challenges in countering adaptive adversarial tactics, requiring intelligent decision-support frameworks. To enhance human-AI collaboration in SOCs, we propose a Cognitive Hierarchy Theory-driven Deep Q-Network (CHT-DQN) framework that models the interactive decision-making process between SOC analysts and AI-driven APT bots. Our approach assumes that the SOC analyst (defender) operates at cognitive level-1, anticipating attacker strategies, while the APT bot (attacker) follows a level-0 policy, making naive exploitative decisions. By incorporating CHT into DQN, the framework enhances adaptive SOC defense strategies using Attack Graph (AG)-based reinforcement learning. Through extensive simulation experiments across varying attack graph complexities, the proposed model consistently achieves higher average weighted data protection and reduced action discrepancies, particularly when SOC analysts utilize CHT-DQN. A theoretical lower bound analysis further validates the superiority of CHT-DQN over standard DQN, demonstrating that as AG complexity increases, CHT-DQN yields higher Q-value performance. A comprehensive human-in-the-loop (HITL) evaluation involving IT security experts was conducted on Amazon Mechanical Turk (MTurk). The findings indicate that SOC analysts who employ CHT-DQN-driven transition probabilities demonstrate greater alignment with adaptive DQN attackers, thereby achieving enhanced data protection. Moreover, human decision patterns in the study align with Prospect Theory, demonstrating risk aversion after failure and risk-seeking behavior after success. The findings highlight the significant potential of integrating cognitive modeling into deep reinforcement learning to improve Security Operations Center functions. This combination presents a promising avenue for developing real-time, adaptive mechanisms for cloud security.
@@ 6.3. Human-Interactive Web-Based DRL Experiments on Amazon MTurk
Figure 8. Node reselection likelihood by human and AI-based SOC analysts, illustrating risk aversion and risk-seeking behaviors.
[…]
@@ 7. Conclusions
This study introduces a Cognitive Hierarchy Theory-driven Deep Reinforcement Learning framework for human-AI collaboration in SOCs, demonstrating its effectiveness in mitigating AI-driven APT threats through adaptive decision-making. The integration of Attack Graphs with CHT-DQN enables SOC analysts to model adversarial behavior at multiple levels of reasoning, significantly outperforming standard DQN across both automated and human-in-the-loop defense scenarios. Key findings from human-AI interaction experiments on Amazon MTurk show that SOC analysts at cognitive level-1 with access to CHT-DQN-driven transition probabilities—incorporating the attacker’s level-0 policy and attack graph likelihoods—achieve higher average weighted data protection ratios and show better alignment with adaptive attackers, as evidenced by reduced action discrepancies. These experiments also exhibit alignment with Prospect Theory’s principles, with SOC analysts showing risk aversion after failure and risk-seeking after success, underscoring the advantages of human-AI synergy in cybersecurity. Simulation experiments further validate that CHT-DQN SOC analysts adapt more effectively to increasing attack graph complexity, demonstrating superior data protection capabilities against both random and strategic attackers. Increasing the AG complexity leads to higher weighted data protection ratios and lower action frequency discrepancies, reinforcing the ability of CHT-DQN to model and counter evolving adversarial strategies.

### v2

@@ Abstract.
Given the complexity of multi-tenant cloud environments and the growing need for real-time threat mitigation, Security Operations Centers (SOCs) must integrate AI-driven adaptive defense mechanisms to counter Advanced Persistent Threats (APTs) effectively. However, SOC analysts face challenges in countering adaptive adversarial tactics, requiring intelligent decision-support frameworks. To enhance human-AI collaboration in SOCs, we propose a Cognitive Hierarchy Theory-driven Deep Q-Network (CHT-DQN) framework that models the interactive decision-making process between SOC analysts and AI-driven APT bots. Our approach assumes that the SOC analyst (defender) operates at cognitive level-1, anticipating attacker strategies, while the APT bot (attacker) follows a level-0 policy, making naive exploitative decisions. By incorporating CHT into DQN, the framework enhances adaptive SOC defense strategies using Attack Graph (AG)-based reinforcement learning. Through extensive simulation experiments across varying attack graph complexities, the proposed model consistently achieves higher average weighted data protection and reduced action discrepancies, particularly when SOC analysts utilize CHT-DQN. A theoretical lower bound analysis further validates the superiority of CHT-DQN over standard DQN, demonstrating that as AG complexity increases, CHT-DQN yields higher Q-value performance. A comprehensive human-in-the-loop (HITL) evaluation involving IT security experts was conducted on Amazon Mechanical Turk (MTurk). The findings indicate that SOC analysts who employ CHT-DQN-driven transition probabilities demonstrate greater alignment with adaptive DQN attackers, thereby achieving enhanced data protection. Moreover, human decision patterns align with Prospect Theory (PT) and Cumulative Prospect Theory (CPT): participants tend to avoid reselection after failure but persist after success. This asymmetry reflects heightened loss sensitivity and biased probability weighting—underestimating gains after loss and overestimating continued success. The findings highlight the significant potential of integrating cognitive modeling into deep reinforcement learning to improve Security Operations Center functions. This combination presents a promising avenue for developing real-time, adaptive mechanisms for cloud security.
@@ 4.3. CHT-DQN Algorithm
Q-value Function of Attacker.
The level-0 agent, which is the cloud attacker, does not use strategic thinking and attempts to reach its goal without considering how the other player might affect its action. For the attacker at level-0:
(13) $Q^{0}_{\mathcal{A}}(s,a_{\mathcal{A}};\theta_{\mathcal{A}})=\sum_{s^{\prime}}p(s^{\prime}|s,a_{\mathcal{A}})\big{[}u_{\mathcal{A}}(s,a_{\mathcal{A}},s^{\prime})+\gamma\max_{a^{\prime}_{\mathcal{A}}}Q^{0}_{\mathcal{A}}(s^{\prime},a^{\prime}_{\mathcal{A}};\theta_{\mathcal{A}})\big{]},$
where $Q^{0}_{\mathcal{A}}$ represents the Q-value for the attacker at level-0, $p(s^{\prime}|s,a_{\mathcal{A}})$ denotes the transition probabilities of the attacker, $u_{\mathcal{A}}$ is the utility (reward) of the attacker, and $\gamma$ is the discount factor.
@@ 6.3. Human-Interactive Web-Based DRL Experiments on Amazon MTurk
Figure 8. Node reselection likelihood by human and AI-based SOC analysts. Human behavior shows reduced reselection after failure and persistent reselection after success. This asymmetry is consistent with the framing effect and probability weighting in Prospect Theory (PT) and Cumulative Prospect Theory (CPT).
Figure 8(a) shows that human SOC analysts in Scenarios 1 and 2 are less likely to reselect nodes that failed to protect data in the previous round, while Figure 8(b) reveals a continued tendency to reselect previously successful nodes. This asymmetry is well explained by Prospect Theory (PT) and Cumulative Prospect Theory (CPT) (Kahneman and Tversky (kahneman1979prospect, ); Tversky and Kahneman (tversky1992advances, )).
According to PT, the subjective value function is concave for gains and convex for losses, with losses perceived more strongly due to loss aversion. This amplifies the psychological cost of failure, discouraging immediate reselection. CPT further refines this perspective through nonlinear probability weighting, where the probability of success after a failure tends to be underweighted, whereas the likelihood of continued success is often overestimated. Together, these mechanisms explain the observed human behavior: avoidance of previously failed nodes and a persistent return to those that previously yielded success.
@@ 7. Conclusions
This study introduces a Cognitive Hierarchy Theory-driven Deep Reinforcement Learning framework for human-AI collaboration in SOCs, demonstrating its effectiveness in mitigating AI-driven APT threats through adaptive decision-making. The integration of Attack Graphs with CHT-DQN enables SOC analysts to model adversarial behavior at multiple levels of reasoning, significantly outperforming standard DQN across both automated and human-in-the-loop defense scenarios. Key findings from human-AI interaction experiments on Amazon MTurk show that SOC analysts at cognitive level-1 with access to CHT-DQN-driven transition probabilities—incorporating the attacker’s level-0 policy and attack graph likelihoods—achieve higher average weighted data protection ratios and show better alignment with adaptive attackers, as evidenced by reduced action discrepancies. These experiments also align with principles from Prospect Theory and Cumulative Prospect Theory, as SOC analysts tend to avoid previously failed actions but persist with those that succeeded—highlighting cognitively grounded patterns in human-AI cybersecurity collaboration. Simulation experiments further validate that CHT-DQN SOC analysts adapt more effectively to increasing attack graph complexity, demonstrating superior data protection capabilities against both random and strategic attackers. Increasing the AG complexity leads to higher weighted data protection ratios and lower action frequency discrepancies, reinforcing the ability of CHT-DQN to model and counter evolving adversarial strategies.